# Finite presentability of twisted Brin–Thompson groups

### v1

@@ 5.1. Boone–Higman via (A)
Every finitely presented, strongly shift-similar group admits an action of type (A), and so in particular has solvable word problem (and satisfies the Boone–Higman conjecture via (A)).
@@
Let $G\leq\operatorname{Sym}(\mathbb{N})$ be finitely presented and strongly shift-similar. We claim the action of $G$ on $\mathbb{N}$ is of type (A). If $G$ is finite there is nothing to prove, so assume $G$ is infinite. By [MZ, Theorem 3.12], we know that $G$ contains $\operatorname{Sym_{fin}}(\mathbb{N})$, the subgroup of elements of $\operatorname{Sym}(\mathbb{N})$ that are the identity outside a finite subset. In particular, the action of $G$ on $\mathbb{N}$ has finitely many orbits of two-element subsets (in fact it is highly transitive). It remains only to prove that each point stabilizer is finitely generated. Let $G_{j}=\operatorname{Stab}_{G}(j)$. Since $G$ is strongly shift-similar, by [MZ, Lemma 3.19] the restriction of $\psi_{j}$ to $G_{j}$ is an isomorphism $\psi_{j}\colon G_{j}\to G$. Hence $G$ is finitely generated (even finitely presented). ∎
@@
We close this subsection by establishing some closure properties for groups satisfying the Boone–Higman conjecture via (A). First we prove closure under direct products.
@@
If two groups satisfy the Boone–Higman conjecture via (A), then so does their direct product.
@@
Let $G$ be a group. If a finite index subgroup of $G$ satisfies the Boone–Higman conjecture via (A), then so does $G$. In particular, satisfying the Boone–Higman conjecture via (A) is a commensurability invariant.
@@
Let $H\leq G$ be a finite index subgroup that satisfies the Boone–Higman conjecture via (A). Say $H$ embeds into a group $E$ admitting an action of type (A) on a set $S$. Up to passing to a deeper finite index subgroup, we may assume that $H$ is normal in $G$. By the Kaloujnine–Krasner extension theorem [KK48], there is an embedding of $G$ into the regular wreath product $H\wr(G/H)$. Now consider the permutational wreath product $E\wr_{n}\Sigma_{n}$ with its natural action on $S\times\{1,\dots,n\}$. Since $\Sigma_{n}$ is finite and the action of $E$ on $S$ is of type (A), it is straightforward to check that the action of $E\wr_{n}\Sigma_{n}$ on $S\times\{1,\dots,n\}$ is of type (A). Choosing $n=|G/H|$, and embedding $G/H$ into $\Sigma_{n}$ via the regular action, we get embeddings
@@
As far as we can tell, neither of these results obviously holds for the Boone–Higman conjecture, without assuming the starting groups satisfy the conjecture via (A). That is, it is not clear how to prove that a direct product of two finitely presented simple groups embeds into a finitely presented simple group, nor how to upgrade a virtual embedding into a finitely presented simple group to a full embedding. Some related combination-type results are unclear to us even for the Boone–Higman conjecture via (A), for example:
@@ 5.2. Cosets and double cosets
Let $G$ be a finitely presented group and $H_{1},\dots,H_{n}\leq G$ finitely generated subgroups. Suppose that the intersection of all conjugates of all the $H_{i}$ is trivial, and that there are finitely many double cosets in $H_{i}\backslash G/H_{j}$, for all $1\leq i,j\leq n$. Then $G$ admits an action of type (A), and hence satisfies the Boone–Higman conjecture via (A).
@@
Let $G=H\bowtie K$ as above, with $G$ finitely presented and $H$ finitely generated. Suppose that the action $\phi$ of $H$ on $K$ is faithful and has finitely many orbits. Then $G$ admits an action of type (A), and hence satisfies the Boone–Higman conjecture via (A).
@@
Let $G$ be a finitely presented group. Suppose that $G$ decomposes as a semidirect product $G=H\ltimes K$ such that the action of $H$ on $K$ is faithful and has finitely many orbits. Then $G$ admits an action of type (A), and hence satisfies the Boone–Higman conjecture via (A).
@@
This is less clearly useful, since now the action of $H$ on $K$ is by automorphisms, and it is rather difficult for a group to have finitely many orbits under actions by automorphisms. Thus, we anticipate the general Zappa–Szép result is more likely to be useful in the future than this semidirect product result.
@@ 5.3. Simple groups themselves
A particularly intriguing question is whether satisfying the Boone–Higman conjecture is equivalent to doing so via (A). In other words, we want to know whether every finitely presented simple group itself embeds into a group admitting an action of type (A). Simple groups make it particularly easy for an intersection of conjugates of a subgroup to be trivial, so we can phrase some interesting sufficient conditions for satisfying the Boone–Higman conjecture via (A).
First recall the Boone–Higman–Thompson theorem, that every finitely generated group with solvable word problem embeds into a finitely generated simple group $H$ that embeds into a finitely presented group $G$ [BH74, Tho80] (so the Boone–Higman conjecture is that this can be accomplished in a single embedding). We now have the following two sufficient conditions for satisfying the Boone–Higman conjecture (via (A)), which are immediate from Proposition 5.8:
@@
Let $H$ be a finitely generated simple subgroup of a finitely presented group $G$, and assume $H$ is not normal in $G$. If there are finitely many double cosets in $H\backslash G/H$, then $G$ admits an action of type (A), and hence satisfies the Boone–Higman conjecture via (A).
@@
Let $G$ be a finitely presented simple group. Suppose there exists a finitely generated proper subgroup $H$ with finitely many double cosets. Then $G$ admits an action of type (A), and hence satisfies the Boone–Higman conjecture via (A).
@@
As a concluding remark, this last result applies to many existing families of finitely presented simple groups, for example it is an easy exercise to find finitely generated proper subgroups with finitely many double cosets inside Thompson’s groups $T$ and $V$, and various groups related to these. It is an interesting problem, however, to try to find finitely generated proper subgroups with finitely many double cosets inside finitely presented simple groups coming from outside the extended family of Thompson-like groups, such as Burger–Mozes groups. We leave this as a question, phrased more generally as follows:

### v2

@@ 5.1. Permutational Boone–Higman
Every finitely presented, strongly shift-similar group admits an action of type (A), and so in particular has solvable word problem (and satisfies the (permutational) Boone–Higman conjecture).
@@
Let $G\leq\operatorname{Sym}(\mathbb{N})$ be finitely presented and strongly shift-similar. We claim the action of $G$ on $\mathbb{N}$ is of type (A). If $G$ is finite there is nothing to prove, so assume $G$ is infinite. By [MZ, Theorem 3.12], we know that $G$ contains $\operatorname{Sym_{fin}}(\mathbb{N})$, the subgroup of elements of $\operatorname{Sym}(\mathbb{N})$ that are the identity outside a finite subset. In particular, the action of $G$ on $\mathbb{N}$ has finitely many orbits of two-element subsets (in fact it is highly transitive). It remains only to prove that each point stabilizer is finitely generated. Let $G_{j}=\operatorname{Stab}_{G}(j)$. Since $G$ is strongly shift-similar, by [MZ, Lemma 3.19] the restriction of $\psi_{j}$ to $G_{j}$ is an isomorphism $\psi_{j}\colon G_{j}\to G$. Hence $G_{j}$ is finitely generated (even finitely presented). ∎
@@
See [MZ, Example 3.31] for more on possible future connections between shift-similar groups and the Boone–Higman conjecture.
We close this subsection by establishing some closure properties for groups satisfying the permutational Boone–Higman conjecture. First we prove closure under direct products.
@@
If two groups satisfy the permutational Boone–Higman conjecture, then so does their direct product.
@@
Let $G$ be a group. If a finite index subgroup of $G$ satisfies the permutational Boone–Higman conjecture, then so does $G$. In particular, satisfying the permutational Boone–Higman conjecture is a commensurability invariant.
@@
Let $H\leq G$ be a finite index subgroup that satisfies the permutational Boone–Higman conjecture. Say $H$ embeds into a group $E$ admitting an action of type (A) on a set $S$. Up to passing to a deeper finite index subgroup, we may assume that $H$ is normal in $G$. By the Kaloujnine–Krasner extension theorem [KK48], there is an embedding of $G$ into the regular wreath product $H\wr(G/H)$. Now consider the permutational wreath product $E\wr_{n}\Sigma_{n}$ with its natural action on $S\times\{1,\dots,n\}$. Since $\Sigma_{n}$ is finite and the action of $E$ on $S$ is of type (A), it is straightforward to check that the action of $E\wr_{n}\Sigma_{n}$ on $S\times\{1,\dots,n\}$ is of type (A). Choosing $n=|G/H|$, and embedding $G/H$ into $\Sigma_{n}$ via the regular action, we get embeddings
@@
As far as we can tell, neither of these results obviously holds for the Boone–Higman conjecture, without assuming the starting groups satisfy the permutational version. That is, it is not clear how to prove that a direct product of two finitely presented simple groups embeds into a finitely presented simple group, nor how to upgrade a virtual embedding into a finitely presented simple group to a full embedding. Some related combination-type results are unclear to us even for the permutational Boone–Higman conjecture, for example:
@@ 5.2. Cosets and double cosets
Let $G$ be a finitely presented group and $H_{1},\dots,H_{n}\leq G$ finitely generated subgroups. Suppose that the intersection of all conjugates of all the $H_{i}$ is trivial, and that there are finitely many double cosets in $H_{i}\backslash G/H_{j}$, for all $1\leq i,j\leq n$. Then $G$ admits an action of type (A), and hence satisfies the (permutational) Boone–Higman conjecture.
@@
Let $G=H\bowtie K$ as above, with $G$ finitely presented and $H$ finitely generated. Suppose that the action $\phi$ of $H$ on $K$ is faithful and has finitely many orbits. Then $G$ admits an action of type (A), and hence satisfies the (permutational) Boone–Higman conjecture.
@@
Let $G$ be a finitely presented group. Suppose that $G$ decomposes as a semidirect product $G=H\ltimes K$ such that the action of $H$ on $K$ is faithful and has finitely many orbits. Then $G$ admits an action of type (A), and hence satisfies the (permutational) Boone–Higman conjecture.
@@
This is less clearly useful, since now the action of $H$ on $K$ is by automorphisms, and it is rather difficult for a group to have finitely many orbits under actions by automorphisms. Thus, we anticipate the general Zappa–Szép result is more likely to be useful in the future than this semidirect product result. We should also mention that, as best we can tell, the solvability of the word problem for a Zappa–Szép product (or even semidirect product) of groups with solvable word problem does not obviously follow, if the actions involved are very wild.
@@ 5.3. Simple groups themselves
A particularly intriguing question is whether satisfying the Boone–Higman conjecture is equivalent to satisfying the permuational version. In other words, we want to know whether every finitely presented simple group itself embeds into a group admitting an action of type (A). Simple groups make it particularly easy for an intersection of conjugates of a subgroup to be trivial, so we can phrase some interesting sufficient conditions for satisfying the permutational Boone–Higman conjecture.
First recall the Boone–Higman–Thompson theorem, that every finitely generated group with solvable word problem embeds into a finitely generated simple group $H$ that embeds into a finitely presented group $G$ [BH74, Tho80] (so the Boone–Higman conjecture is that this can be accomplished in a single embedding). We now have the following two sufficient conditions for satisfying the (permutational) Boone–Higman conjecture, which are immediate from Proposition 5.8:
@@
Let $H$ be a finitely generated simple subgroup of a finitely presented group $G$, and assume $H$ is not normal in $G$. If there are finitely many double cosets in $H\backslash G/H$, then $G$ admits an action of type (A), and hence satisfies the (permutational) Boone–Higman conjecture.
@@
Let $G$ be a finitely presented simple group. Suppose there exists a finitely generated proper subgroup $H$ with finitely many double cosets. Then $G$ admits an action of type (A), and hence satisfies the permutational Boone–Higman conjecture.
@@
As a concluding remark, this last result applies to many existing families of finitely presented simple groups, for example it is an easy exercise to find finitely generated proper subgroups with finitely many double cosets inside Thompson’s groups $T$ and $V$, and various groups related to these. It is an interesting problem, however, to try to find finitely generated proper subgroups with finitely many double cosets inside finitely presented simple groups coming from outside the extended family of Thompson-like groups. We leave this as a question, phrased more generally as follows:
@@
  2. (ii)
    Does every finitely presented simple group embed in a group admitting an action of type (A)? (And hence satisfy the permutational Boone–Higman conjecture?)
We should mention that for Burger–Mozes groups, which are outside the Thompson group family, this second question was proved to be true in [BLIW], and so Burger–Mozes groups satisfy the permutational Boone–Higman conjecture. Thus, essentially the only known finitely presented simple groups outside the Thompson family for which the permutational Boone–Higman conjecture is open are the non-affine Kac–Moody groups used by Caprace and Rémy in [CR09, CR10].
To be clear, if the answer to Question 5.13(ii) is always “yes”, then the permutational and regular Boone–Higman conjectures are equivalent, and finitely presented twisted Brin–Thompson groups are in some sense universal among finitely presented simple groups.